OpenStudy (anonymous):

Tan(tan^-1 7/3) Find the exact value

OpenStudy (anonymous):

my guess is \(\frac{7}{3}\) since this is like "who is buried in grant's tomb" you are composing a function with its inverse you always get the input back

OpenStudy (anonymous):

Another example is Sin^-1 (sin11pi/5) @satellite73

OpenStudy (anonymous):

now that is a different story

OpenStudy (anonymous):

the answer CANNOT be \(\frac{11\pi}{5}\) because the range of arcsine is \([-\frac{\pi}{2},\frac{\pi}{2}]\)

OpenStudy (anonymous):

subtract \(2\pi\) from your input to give you the correct answer

OpenStudy (rational):

For trig functions I think below is always true \[f(f^{-1}(x)) = x\] But below wont work if \(x\) is not in the domain of inverse \[f^{-1}(f(x)) \stackrel{*}{=} x\]

OpenStudy (anonymous):

How much would I be subtracting? I don't how much 2pi would take away